# Consumption Partial Insurance

### v5

@@ 19 ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{3})$
20 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}c_{it})^{2})$
21 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}c_{it})$
22 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{3})$
23 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$
24 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times(\Delta^{2}c_{it})^{2})$
25 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}c_{it-2})$
26 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it-2})^{3})$
Notes: The table lists the moments targeted in the estimation of the different specifications of the consumption process. In all cases, the notation $\Delta^{2}x_{t}=\Delta x_{t}+\Delta x_{t-1}=x_{t}-x_{t-2}$ indicates the first difference of the generic variable $x$ between periods $t$ and $t-2$.
@@ B.4.2 Moments of consumption – linear function
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ $=2\sigma_{\xi}^{2}+2(\phi^{(1)})^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\sigma_{v}^{2}+2\sigma_{u^{c}}^{2}$
@@
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\sigma_{\zeta}^{2}+\psi^{(1)}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3})$ $=2\gamma_{\xi}+2(\phi^{(1)})^{3}\gamma_{\zeta}+(\psi^{(1)})^{3}\gamma_{v}$
@@
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\gamma_{\zeta}+\psi^{(1)}\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{2})$ $=2(\phi^{(1)})^{2}\gamma_{\zeta}+(\psi^{(1)})^{2}\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$ $=\psi^{(1)}\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it-2})^{2})$ $=-(\psi^{(1)})^{2}\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$ $=0$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$ $=-\psi^{(1)}\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$ $=0$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$ $=0$
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{4})$ $=2\kappa_{\xi}+2(\phi^{(1)})^{4}\kappa_{\zeta}+(\psi^{(1)})^{4}\kappa_{v}+2\kappa_{u^{c}}+6(\sigma_{\xi}^{2})^{2}+6(\phi^{(1)})^{4}(\sigma_{\zeta}^{2})^{2}+6(\sigma_{u^{c}}^{2})^{2}$
$+24(\phi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+12(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+24\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}$
$+12(\phi^{(1)})^{2}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+24(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+12(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=4(\sigma_{\xi}^{2})^{2}+8(\phi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+4(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+8\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}+4(\phi^{(1)})^{4}(\sigma_{\zeta}^{2})^{2}+3(\sigma_{u^{c}}^{2})^{2}$
$+4(\phi^{(1)})^{2}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+8(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+(\psi^{(1)})^{4}(\sigma_{v}^{2})^{2}+4(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}+\kappa_{u^{c}}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3}\times\Delta^{2}c_{it-2})$ $=-\kappa_{u^{c}}-3(\sigma_{u^{c}}^{2})^{2}-6\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}-6(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}-3(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{3})$ $=-\kappa_{u^{c}}-3(\sigma_{u^{c}}^{2})^{2}-6\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}-6(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}-3(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}c_{it})^{2})$ $=4\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+2(\phi^{(1)})^{2}\kappa_{\zeta}+6(\phi^{(1)})^{2}(\sigma_{\zeta}^{2})^{2}+2(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+4\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}$
$+4\sigma_{v}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{v}^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\kappa_{v}+(\psi^{(1)})^{2}(\sigma_{v}^{2})^{2}+4\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
$+4\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}+2(\psi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{v}^{2}+4\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}+8\phi^{(1)}\psi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\kappa_{\zeta}+6\phi^{(1)}(\sigma_{\zeta}^{2})^{2}+\psi^{(1)}\kappa_{v}+3\psi^{(1)}(\sigma_{v}^{2})^{2}$
$+6\psi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}+12\phi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}+12\phi^{(1)}\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}+6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{3})$ $=2(\phi^{(1)})^{3}\kappa_{\zeta}+6(\phi^{(1)})^{3}(\sigma_{\zeta}^{2})^{2}+(\psi^{(1)})^{3}\kappa_{v}+12\phi^{(1)}\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+6\psi^{(1)}\sigma_{\xi}^{2}\sigma_{v}^{2}$
$+6\phi^{(1)}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+6(\phi^{(1)})^{2}\psi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}+12\phi^{(1)}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=4\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}(\sigma_{\zeta}^{2})^{2}+2(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+4\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}$
$+4\sigma_{v}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{v}^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\kappa_{v}+(\psi^{(1)})^{2}(\sigma_{v}^{2})^{2}+4\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
$+4\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}+2(\psi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{v}^{2}+4\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times(\Delta^{2}c_{it})^{2})$ $=4\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}(\sigma_{\zeta}^{2})^{2}+2(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+4\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}$
$+4\sigma_{v}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{v}^{2}\sigma_{\zeta}^{2}+2(\psi^{(1)})^{2}(\sigma_{v}^{2})^{2}+4\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
$+4\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}+2(\psi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{v}^{2}+4\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\kappa_{v}-3\psi^{(1)}(\sigma_{v}^{2})^{2}-6\psi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}-6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it-2})^{3})$ $=-(\psi^{(1)})^{3}\kappa_{v}-6\psi^{(1)}\sigma_{v}^{2}\sigma_{\xi}^{2}-6(\phi^{(1)})^{2}\psi^{(1)}\sigma_{v}^{2}\sigma_{\zeta}^{2}-6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$

### v6

@@ 19 ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{3})$
20 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}c_{it})^{2})$
21 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3}\times\Delta^{2}c_{it})$
22 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it})^{3})$
23 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$
24 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times(\Delta^{2}c_{it})^{2})$
25 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3}\times\Delta^{2}c_{it-2})$
26 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it-2})^{3})$
Notes: The table lists the moments targeted in the estimation of the different specifications of the consumption process. The notation $\Delta^{2}x_{t}=x_{t}-x_{t-2}$ indicates a first difference of the generic variable $x$ between periods $t$ and t2t-2; $\Delta^{2}_{\rho}x_{t}=x_{t}-\rho^{2}x_{t-2}$ indicates the analogous quasi-difference.
@@ B.4.2 Moments of consumption – linear function
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ $=2\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\sigma_{v}^{2}+2\sigma_{u^{c}}^{2}$
@@
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it})$ $=\phi^{(1)}(1+\rho^{2})\sigma_{\zeta}^{2}+\psi^{(1)}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\rho^{2}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3})$ $=2\gamma_{\xi}+(\phi^{(1)})^{3}(1+\rho^{3})\gamma_{\zeta}+(\psi^{(1)})^{3}\gamma_{v}$
@@
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it})$ $=\phi^{(1)}(1+\rho^{3})\gamma_{\zeta}+\psi^{(1)}\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it})^{2})$ $=(\phi^{(1)})^{2}(1+\rho^{3})\gamma_{\zeta}+(\psi^{(1)})^{2}\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it-2})$ $=\psi^{(1)}\rho^{4}\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it-2})^{2})$ $=-(\psi^{(1)})^{2}\rho^{2}\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times\Delta^{2}c_{it})$ $=0$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it})$ $=-\psi^{(1)}\rho^{2}\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it-2})$ $=0$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it+2})$ $=0$
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{4})$ $=2\kappa_{\xi}+(\phi^{(1)})^{4}(1+\rho^{4})\kappa_{\zeta}+(\psi^{(1)})^{4}\kappa_{v}+2\kappa_{u^{c}}+6(\sigma_{\xi}^{2})^{2}$
$+6(\phi^{(1)})^{4}\rho^{2}(\sigma_{\zeta}^{2})^{2}+6(\sigma_{u^{c}}^{2})^{2}+12(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\xi}^{2}\sigma_{\zeta}^{2}$
$+12(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+24\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}+6(\phi^{(1)})^{2}(\psi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
$+12(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+12(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=4(\sigma_{\xi}^{2})^{2}+4(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+4(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+8\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}$
$+(\phi^{(1)})^{4}(1+\rho^{2})^{2}(\sigma_{\zeta}^{2})^{2}+3(\sigma_{u^{c}}^{2})^{2}+2(\phi^{(1)})^{2}(\psi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
$+4(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+(\psi^{(1)})^{4}(\sigma_{v}^{2})^{2}+4(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}+\kappa_{u^{c}}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3}\times\Delta^{2}c_{it-2})$ $=-\kappa_{u^{c}}-3(\sigma_{u^{c}}^{2})^{2}-6\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}-3(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}-3(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{3})$ $=-\kappa_{u^{c}}-3(\sigma_{u^{c}}^{2})^{2}-6\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}-3(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}-3(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}c_{it})^{2})$ $=2(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{4})\kappa_{\zeta}+6(\phi^{(1)})^{2}\rho^{2}(\sigma_{\zeta}^{2})^{2}$
$+(\psi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+2(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+2(1+\rho^{4})\sigma_{v}^{2}\sigma_{\xi}^{2}$
$+(\phi^{(1)})^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{v}^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\kappa_{v}+(\psi^{(1)})^{2}\rho^{4}(\sigma_{v}^{2})^{2}$
$+2(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{c}}^{2}+2(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}$
$+(\psi^{(1)})^{2}(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{v}^{2}+2(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}+4\phi^{(1)}\psi^{(1)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3}\times\Delta^{2}c_{it})$ $=\phi^{(1)}(1+\rho^{4})\kappa_{\zeta}+6\phi^{(1)}\rho^{2}(\sigma_{\zeta}^{2})^{2}+\psi^{(1)}\kappa_{v}+3\psi^{(1)}\rho^{4}(\sigma_{v}^{2})^{2}$
$+3\psi^{(1)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+3\phi^{(1)}(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
$+3\phi^{(1)}(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}+3\psi^{(1)}(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it})^{3})$ $=(\phi^{(1)})^{3}(1+\rho^{4})\kappa_{\zeta}+6(\phi^{(1)})^{3}\rho^{2}(\sigma_{\zeta}^{2})^{2}+(\psi^{(1)})^{3}\kappa_{v}$
$+6\phi^{(1)}(1+\rho^{2})\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+6\psi^{(1)}\sigma_{\xi}^{2}\sigma_{v}^{2}+3\phi^{(1)}(\psi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
$+3(\phi^{(1)})^{2}\psi^{(1)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+6\phi^{(1)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=2(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})^{2}(\sigma_{\zeta}^{2})^{2}+(\psi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
$+2(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+2(1+\rho^{4})\sigma_{v}^{2}\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{v}^{2}\sigma_{\zeta}^{2}$
$+(\psi^{(1)})^{2}\rho^{4}\kappa_{v}+(\psi^{(1)})^{2}(\sigma_{v}^{2})^{2}+2(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{c}}^{2}+2(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}$
$+(\phi^{(1)})^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{v}^{2}+2(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times(\Delta^{2}c_{it})^{2})$ $=2(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})^{2}(\sigma_{\zeta}^{2})^{2}+(\psi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
$+2(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+2(1+\rho^{4})\sigma_{v}^{2}\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{v}^{2}\sigma_{\zeta}^{2}$
$+(\psi^{(1)})^{2}(1+\rho^{4})(\sigma_{v}^{2})^{2}+2(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{c}}^{2}+2(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}$
$+(\phi^{(1)})^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{v}^{2}+2(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\rho^{6}\kappa_{v}-3\psi^{(1)}\rho^{2}(\sigma_{v}^{2})^{2}-3\psi^{(1)}(1+\rho^{2})\rho^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}-3\psi^{(1)}(1+\rho^{4})\rho^{2}\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it-2})^{3})$ $=-(\psi^{(1)})^{3}\rho^{2}\kappa_{v}-6\psi^{(1)}\rho^{2}\sigma_{v}^{2}\sigma_{\xi}^{2}-3(\phi^{(1)})^{2}\psi^{(1)}(1+\rho^{2})\rho^{2}\sigma_{v}^{2}\sigma_{\zeta}^{2}-6\psi^{(1)}\rho^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$